TWO-ROTOR CALCULATIONS FOR MOLECULES WITH C_{2} OR C_{s} SYMMETRY

\underline{PETER GRONER}, Department of Chemistry, University of Missouri - Kansas City, Kansas City, MO 64110.

An effective rotation-internal motion Hamiltonian for molecules with two periodic internal motions ^{a} has been applied recently to dimethyl ether ^{b} and acetone ^{c} whose equilibrium structures have C_{2v} symmetry. The program used to analyze their rotational spectra has been modified to allow treatment of two-rotor molecules with lower symmetry (equivalent rotor problems for molecules with C_{s} and C_{2} symmetry and nonequivalent rotor problems with and without a plane of symmetry). It has been used to analyze the microwave literature data of CH_{3}SSCH_{3} (C_{2} symmetry, ground state) ^{d}, CH_{3}CHFCH_{3} (C_{s} symmetry, ground and two excited states) ^{e}, and CH_{3}OSiH_{3} (nonequivalent, C_{s} symmetry, ground state) ^{f}. The experimental data have been fitted to experimental precision by the nonlinear least-squares method to determine spectroscopic parameters (rotational constants, centrifugal distortion constants and tunneling parameters). Applications to other molecules are in progress.

aP. Groner, J. Chem. Phys. 107, 4483 (1997)

bP. Groner, S. Albert, E. Herbst and F. C. De Lucia, Astrophys. J. 500, 1059 (1998)

cP. Groner, S. Albert, E. Herbst, F. C. De Lucia, F. J. Lovas, J. C. Pearson and B. J. Drouin, Astrophys. J. (submitted)

dM. Meyer, J. Mol. Struct. 273, 99 (1992)

eM. Meyer and H. Dreizler, Z. Naturforsch. A 43, 138 (1988); J. Mol. Spectrosc. 148, 310 (1991)

fC. D. LeCroix, R. F. Curl, P. M. McKinney and R. J. Myers, J. Mol. Spectrosc. 53, 250 (1972)